### 11 Mar 2026 v1

# A note on Ramsey numbers for minors

Abstract
Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of a complete graph on $n$ vertices in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$, i.e., a graph that could be transformed into a clique on $k$ vertices via a sequence of edge contractions and vertex deletions. We show that
$k\sqrt{\log k}(1+o(1))\leq R_{h}(k;2)\leq 1.031\cdot k\sqrt{\log k}(1+o(1))$
and for a constant $\beta=0.265656$...,
$R_{h}(k;\ell)=2\beta\ell k\sqrt{\log k}(1+o(1)).$
1 Introduction
The classical Ramsey number $R(k)$ is defined as the smallest integer $n$ such that any edge coloring of a complete graph $K_{n}$ on $n$ vertices with two colors contains a monochromatic $K_{k}$. This notion was generalised in many different ways, in particular “through parameter”. Specifically, when $\xi$ is some graph parameter, then $R_{\xi}(k)$ is defined as the smallest $n$ such that any edge coloring of $K_{n}$ in two colors contains a monochromatic subgraph $G$ with $\xi(G)=k$. So, $R(k)=R_{\omega}(k)$, where $\omega(G)$ is the clique number of $G$. One of the classical graph parameters is Hadwiger number $h(G)$ that is the largest $k$ so that $G$ contains a $K_{k}$-minor, that means that $K_{k}$ can be obtained from $G$ by edge-contractions and vertex-deletions. In this note, we shall discuss the Ramsey number for minors $R_{h}(k)$. Although it is very natural and complements extensive research on Hadwiger conjecture, it seems that it never was explicitly addressed in the literature. The Hadwiger conjecture [6] states that for any graph $G$, $h(G)\geq\chi(G)$, where $\chi(G)$ is the chromatic number of $G$, see for example Delcourt and Postle [5] and Hendrey, Norin, Steiner, and Turcotte [7]. A related density question is to bound $h(G)$ in terms of the average degree of $G$. For each $k\geq 2$, let
$c(k):={\rm min}\{c:|E(G)|\geq c|V(G)|\Rightarrow h(G)\geq k\}.$
The fact that one could take minimum and not just an infimum in the above definition is due to Mader [9]. Thomason [13] determined $c(k)$ asymptotically, see some related papers [1, 14, 15, 16]. Let $\lambda<1$ be the solution of the equation $1-\lambda+2\lambda\ln\lambda=0$ and and $\beta=(1-\lambda)/2\sqrt{\ln(1/\lambda)}/\sqrt{\log_{2}e}$. Here, $\beta=0.3190863.../\sqrt{\log_{2}e}=0.265656$....
Theorem 1.1 (Thomason [13]).
We have
$c(k)=\beta k\sqrt{\log_{2}k}(1+o(1)).$
The lower bound on $c(k)$ in the above theorem is achieved by the Erdős-Renyi random graph $G(n,p)$, whose Hadwiger number is well understood.
Theorem 1.2 (Bollobás, Catlin, and Erdős [4]).
For a fixed $p$ and q=1pq=1-p,
$h(G(n,p))=(1+o(1))n\left(\frac{\log(1/q)}{\log n}\right)^{1/2}.$
In addition, Thomason [13] proved that the Hadwider number for graphs whose connectivity is not too small is at least one of a random graph with the same density.
Theorem 1.3 (Theorem 4.1, Thomason [13]).
Let $0<\varepsilon<1$. There is $n_{0}=n_{0}(\varepsilon)$, such that if $G$ is a graph on $n$ vertices, $n>n_{0}$, of density $p$, and connectivity at least $n(\log\log\log n)/(\log\log n)$, then
$h(G)\geq\left\lceil(1-\varepsilon)n\left(\frac{\log(1/q^{*})}{\log n}\right)^{1/2}\right\rceil,$
where $q^{*}=\max\{1-p,(\log n)^{-1/\varepsilon}\}$.
Kostochka [8], Stiebitz [12], as well as Reed and Thomas [11] studied Hadwiger number for graphs and their complements in an “off-diagonal” setting when $h(G)$ is a fixed constant. For the rest on this note, the logarithms are base $2$, i.e., $\log=\log_{2}$ and $\beta$ is a constant defined above. We drop floors and ceilings where clear from context.
Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of $K_{n}$ in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$. The following theorems bring known results together to get the bounds on $R_{h}(k;\ell)$. Recall that $R_{h}(k)=R_{h}(k;2)$.
Theorem 1.4.
We have
$k\sqrt{\log k}(1+o(1))\leq R_{h}(k)\leq 2\sqrt{\beta}k\sqrt{\log k}(1+o(1))\leq 1.031\cdot k\sqrt{\log k}(1+o(1)).$
Moreover, $R_{h}(3)=5$, $R_{h}(4)=7$, and for any $k\geq 5$, $2(k-2)\leq R_{h}(k)\leq 32(k-2)\lfloor\log(k-2)\rfloor+1.$
Theorem 1.5.
Let $\beta=0.2656568$... as above. For any $\ell\geq 1$, $R_{h}(3;\ell)=2\ell+1$. For any $\varepsilon>0$ there is $k_{0}=k_{0}(\varepsilon)$, such that for any $k\geq k_{0}$ there is $\ell_{0}=\ell_{0}(k,\varepsilon)$, so that for any $\ell>\ell_{0}$,
$2\beta\ell k\sqrt{\log k}(1-\epsilon)\leq R_{h}(k;\ell)\leq 2\beta\ell k\sqrt{\log k}(1+\epsilon).$
2 Proofs of the main results
First, we note that we can get an upper bound $R_{h}(k)\leq 4\beta k\sqrt{\log k}(1+o(1))\approx 1.0623k\sqrt{\log k}(1+o(1))$ from Theorem 1.1 immediately. Indeed, consider an arbitrary edge coloring of $K_{n}$ with two colors, where $n=4\beta k\sqrt{\log k}(1+o(1))$. Then the majority color class $G$ has size at least $\frac{n^{2}}{4}(1+o(1))=\beta k\sqrt{\log k}(1+o(1))n.$ From Theorem [13], we see that $h(G)\geq k$.
Now, we shall give a more refined argument to improve the constant term from $1.0623$... to $1.03108$.... Let $\varepsilon>0$ be a small constant and $n_{0}=n_{0}(\varepsilon)$ be sufficiently large, in particular satisfying assumptions of Theorem 1.3 and such that $(\log n)^{-1/\varepsilon}$ is sufficiently small, say at most $0.001$. Let $c=2\sqrt{\beta}+\varepsilon$. Note that for a small $\varepsilon$, $c>1/(1-2\varepsilon)$. Consider an arbitrary edge coloring of $K_{n}$ in red and blue, where $n=ck\sqrt{\log k}(1+o(1))$, and $k$ is sufficiently large so that $n>n_{0}$. Let without loss of generality red be a majority color class and $G$ be the graph whose edge set consists of all red edges. We define a function $x(n):=n(\log\log\log n)/(\log\log n)$.
Case 1. Graph $G$ has connectivity at least $x(n)$.
Let $p$ be the density of $G$. Note that $p$ and 1p1-p are larger than $(\log n)^{-1/\varepsilon}$, otherwise we can apply Theorem 1.1 to the red or the blue graph right away and get a monochromatic $K_{k}$-minor. Using the connectivity assumption and Theorem 1.3 with $q^{*}=(1-p)\leq 1/2$, we have $h(G)\geq(1-\varepsilon)n\sqrt{\log(1/2)/\log n}\geq ck(1-2\varepsilon)\geq k$.
Case 2. There is an induced subgraph $G^{\prime}$ of $G$ on at least $n^{\prime}=c^{\prime}n$ vertices, where $cc^{\prime}\geq 1/(1-3\varepsilon)$, minimum degree at least $x(n^{\prime})+k+1$, and density at least $1/2$.
If the connectivity of $G^{\prime}$ is at least $x(n^{\prime})$, then again by Theorem 1.3, we have that $h(G^{\prime})\geq(1-\varepsilon)c^{\prime}n/\sqrt{\log c^{\prime}n}\geq(1-2\varepsilon)cc^{\prime}k\sqrt{\log k}/\sqrt{\log k(1+o(1))}\geq k$. If the connectivity of $G^{\prime}$ is less than $x(n^{\prime})$, consider a smallest vertex cut $X$. By the condition on the minimum degree we see that any connected component of $G^{\prime}-X$ has size at least $k$. Thus there is a $K_{k,k}$ in the complement of $G$, i.e., a blue $K_{k}$-minor.
Consider a largest set $\{v_{1},\ldots,v_{y}\}$ of $y$ vertices such that $deg(v_{1})$ in $G$ and $deg(v_{i})$ in $G-\{v_{1},\ldots,v_{i-1}\}$ is at most $x(n)+k$, for any $2\leq i\leq y$. Let $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$. Then
$\frac{|E(G^{\prime})|}{|V(G^{\prime})|}\geq\frac{|E(G)|-y(x(n)+k)}{|V(G)|-y}\geq\frac{n/4-yx(n)/n-k/n}{1-y/n}.$
In particular, we have $|E(G^{\prime})|/|V(G^{\prime})|\geq|V(G^{\prime})|/4$ for any $y\geq 1$. If $y\geq n(1-c/(4\beta)+\varepsilon)$, then $|E(G^{\prime})|/|V(G^{\prime})|\geq(\beta+\varepsilon)k\sqrt{\log k}$. In this case by Theorem 1.1, $h(G)\geq k$. If $y<n(1-c/(4\beta)+\varepsilon)$, then $|V(G^{\prime})|>n(c/(4\beta)-\varepsilon)$. Thus, we see $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$ satisfies the conditions of Case 2 since $G^{\prime}$ has $c^{\prime}n$ vertices for $c^{\prime}\geq c/(4\beta)-\varepsilon$ and for sufficiently small $\varepsilon$ we have $cc^{\prime}\geq c^{2}/4\beta-c\varepsilon=1+\varepsilon/\sqrt{\beta}+\varepsilon^{2}/(4\beta)-2\sqrt{\beta}\varepsilon-\varepsilon^{2}>1/(1-3\varepsilon)$. This proves the upper bound.
For the lower bound we consider a random coloring of $K_{n}$, $n=k\sqrt{\log k}(1-\epsilon)$, where each edge is colored red with probability $1/2$ and blue with probability $1/2$. Each color class corresponds to $G(n,1/2)$. By Theorem 1.2, almost surely $h(G(n,1/2))<k$. This proves the first statement of the theorem.
Now we prove the bounds for small $k$ and give general, but not optimal bounds that hold for any $k$. To see the lower bound of 2(k2)2(k-2), consider a union of two red vertex-disjoint cliques on k2k-2 vertices each so that all other edges are blue. The upper bound follows from Theorem 1.16 by Bollobás [3], that claims that $c(k)\leq 8(k-2)\lfloor\log_{2}(k-2)\rfloor$. Indeed, let $G$ be a majority color class in a given red/blue coloring of the edges of $K_{n}$. It has at least n(n1)/4n(n-1)/4 edges. Since $(n-1)/4\geq 8(k-2)\lfloor\log_{2}(k-2)\rfloor$, we have that $h(G)\geq k$.
Finally, we treat the small cases $k\in\{3,4\}$. To see that $R_{h}(3)=5$, consider an arbitrary edge coloring of $K_{5}$ in two colors. A largest color class has at least $5$ edges and thus has a cycle, that is a $K_{3}$-minor. On the other hand, a coloring of $K_{4}$ where each color class is isomorphic to $P_{4}$ contains no cycles and thus no $K_{3}$-minors.
For $R_{h}(4)$, we shall use the fact that the largest number of edges in an $n$-vertex graph with no $K_{4}$-minor is at most 2n32n-3, for $n\geq 3$, see for example Bollobás [3]. To see that $R_{h}(4)>6$, consider a 6-cycle and three additional new edges forming a triangle, color this graph red and its complement blue. Since any $K_{4}$-minor has at least four vertices of degree at least three, we see that the color classes are $K_{4}$-minor free. We can immediately see that $R_{h}(4)\leq 8$. Indeed, in any 2-edge coloring of $K_{8}$, a majority color class has size at least $14>2\cdot 8-3$ and thus contains a $K_{4}$-minor. Now, we shall show a stronger bound $R_{h}(4)\leq 7$. Consider an arbitrary red/blue edge coloring of $K_{7}$ with no monochromatic $K_{4}$-minor and let $G$ be a graph whose edge set consists of the red edges. We make several observations. As above $10\leq|E(G)|\leq 11$. Both $G$ and $\overline{G}$ are not bipartite since otherwise one of the parts has size at least $4$. By a special case of Hadwiger conjecture, $\chi(G)=\chi(\overline{G})=3$. Thus $G$ and $\overline{G}$ are three-partite graphs with parts of sizes at most three, i.e., of sizes $3,3,1$ or $3,2,2$. There is no monochromatic $K_{3,3}$ without an edge. If one color class, say blue, has two vertex disjoint triangles, then there are two blue edges between them or three edges forming a $P_{4}$ between them. By considering the remaining vertex, one can see that there is a monochromatic copy of $K_{4}$-minor, a contradiction. Both the red and the blue graphs are three-chromatic with color classes of sizes $3,2,2$. This leaves a final not so long case analysis to demonstrate the existence of a monochromatic $K_{4}$-minor, a contradiction. ∎
To see that $R_{h}(3;\ell)\geq 2\ell+1$, consider a decomposition of $K_{2\ell}$ into $\ell$ Hamiltonian paths. Since any $K_{3}$-minor contains a cycle, we have that this coloring has no monochromatic $K_{3}$-minors. To see that $R_{h}(3;\ell)\leq 2\ell+1$, consider an arbitrary edge-coloring of $K_{2\ell+1}$ in $\ell$ colors. Then some color class has at least $2\ell+1$ edges and thus contains a cycle, that is a $K_{3}$-minor.
The upper bound of the theorem holds in a stronger form, for any $\ell$. Let $n=R_{h}(k;\ell)-1$ and consider an edge-coloring of $K_{n}$ in $\ell$ colors with no monochromatic $K_{k}$-minor. From Theorem 1.1, we know that any $K_{k}$-minor free graph on $n$ vertices has at most $\beta k\sqrt{\log k}n(1+o_{k}(1))$ edges. Since the total number of edges on $n$ vertices is $\binom{n}{2}$, we have that $\binom{n}{2}\leq\ell\cdot\beta k\sqrt{\log k}n(1+o_{k}(1))$. This implies that $n\leq\ell\cdot 2\beta k\sqrt{\log k}(1+o_{k}(1))$.
For the lower bound, let $\varepsilon$ be a small constant, $0<\varepsilon<1$, $k$ be sufficiently large. Consider a graph $H^{\prime}$ on $k\sqrt{\log k}$ vertices and $m=\beta k^{2}\log k(1-\varepsilon^{2}/5)$ edges that is $K_{k}$-minor free, that exists by Theorem 1.1. Let $H$ be obtained from $H^{\prime}$ by attaching a pendant edge. Wilson [17] showed that $K_{n}$ is a pairwise edge-disjoint union of copies of $H$ if $n$ is sufficiently large, $|E(H)|$ divides $\binom{n}{2}$, and the greatest common divisor of the degrees of $H$ divides n1n-1. Since there is a vertex of degree $1$ in $H$, the last condition is satisfied for any $n\geq 2$. Let $\ell$ be sufficiently large, so that for some $n$ satisfying $\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon)\leq n\leq\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon/2)$, $n$ satisfies conditions of Wilson’s theorem, in particulst $|E(H)|$ divides $\binom{n}{2}$. Then $H$ decomposes $K_{n}$. Alon and Yuster [2] proved that if a fixed graph $H$ decomposes $K_{n}$, then there is a coloring of $E(K_{n})$ in at most (1+o(1))(n1)|V(H)|/(2|E(H)|)(1+o(1))(n-1)|V(H)|/(2|E(H)|) colors, such that each color class is a pairwise vertex-disjoint union of copies of $H$. So, we take $\ell$ sufficiently large that $o(1)$ term is at most $\varepsilon/2$. That is, every color class is almost an $H$-factor and it is $K_{k}$-minor free. Since the number of colors is at most $(1+\varepsilon/2)(n-1)|V(H)|/(2|E(H)|)\leq(1+\varepsilon/2)\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon/2)/(2\beta k\sqrt{\log k}(1-\varepsilon^{2}/5))\leq\ell$, we have a coloring of $E(K_{n})$ in $\ell$ colors with no monochromatic copy of a $K_{k}$-minor. ∎
3 Concluding remarks
We showed that $R_{h}(k;2)=ck\sqrt{\log k}(1+o(1))$, where $c$ is a constant, $1\leq c\leq 1.03$.... We believe that $c=1$. It would be interesting to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
To prove the lower bound on $R_{h}(k;\ell)$, one can also use the following packing result by taking $G$ being an $H$-factor, where $H$ is a graph on $k\sqrt{\log k}(1+o(1))$ vertices, density $\beta k\sqrt{\log k}$ and no $K_{k}$-minor. Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ own color. Decompose the remaining edges of $K_{n}$ into forests using Nash-Williams theorem and give each such forest a new own color.
Theorem 3.1 (Messuti, Rödl, Schacht [10]).
For any $\varepsilon>0$ and any $\Delta\in\mathbb{N}$, for any non-trivial minor closed family $\mathcal{G}$ there is $n_{0}$ such that for any $n\geq n_{0}$ and any integer $\ell$ the following holds. If $G\in\mathcal{G}$, $\Delta(G)\leq\Delta$ and $|V(G)|\leq n$, and $\ell|E(G)|\leq\binom{n}{2}$ then $\ell$ copies of $G$ pack in $K_{(1+\varepsilon)n}$.
4 Acknowledgements
The author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.